A cylindrical tank has an 18 inch diameter and is 20 inches high. If it is half full of water, how much more water will be needed to raise it to a height of 18 inches? If one cubic foot of water is about 7.5 gallons of water, how many gallons of water are now in the tank? If one cubic foot of water weight about 63 pounds, how much will the water in the tank weigh?
Answer by lwsshak3(11628)

A cylindrical tank has an 18 inch diameter and is 20 inches high. If it is half full of water, how much more water will be needed to raise it to a height of 18 inches? If one cubic foot of water is about 7.5 gallons of water, how many gallons of water are now in the tank? If one cubic foot of water weight about 63 pounds, how much will the water in the tank weigh?
..
volume of tank=Area*height = pi*r^2*h=3.14*(18/2)^2*20
=3.14*9^2*20=5089.38 in^3
convert to ft^3
ft^3=144 in^3
5089.38/144=35.34 ft^3
1/2 volume of tank=17.67 ft^3
gallons of water in half-full tank=17.67*7.5=132.54 gallons
Weight of water in half-full tank=17.67*63=1113.21 pounds